Ask your own question, for FREE!
Mathematics 9 Online
OpenStudy (anonymous):

if the probability of seeing a moose in a day in a certain region of alaska is 0.39, what is the probability of not seeing any moose in a day there?

OpenStudy (anonymous):

Any guesses?

OpenStudy (kropot72):

The two events 'see a moose in a day' and 'not see a moose in a day' are mutually exclusive. P(A U B) = P(A) + P(B) The probability of seeing a moose in a day or not seeing a moose in a day is 1. Let the probability of not seeing a moose in a day = P(n) 1 = 0.39 + P(n) P(n) = 1 - 0.39

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!